# Kernel PCA for Out-of-Distribution Detection

### v1

@@ Abstract
Out-of-Distribution (OoD) detection is vital for the reliability of Deep Neural Networks (DNNs). Existing works have shown the insufficiency of Principal Component Analysis (PCA) straightforwardly applied on the features of DNNs in detecting OoD data from In-Distribution (InD) data. The failure of PCA suggests that the network features residing in OoD and InD are not well separated by simply proceeding in a linear subspace, which instead can be resolved through proper nonlinear mappings. In this work, we leverage the framework of Kernel PCA (KPCA) for OoD detection, seeking subspaces where OoD and InD features are allocated with significantly different patterns. We devise two feature mappings that induce non-linear kernels in KPCA to advocate the separability between InD and OoD data in the subspace spanned by the principal components. Given any test sample, the reconstruction error in such subspace is then used to efficiently obtain the detection result with $\mathcal{O}(1)$ time complexity in inference. Extensive empirical results on multiple OoD data sets and network structures verify the superiority of our KPCA-based detector in efficiency and efficacy with state-of-the-art OoD detection performances.
@@ 1 Introduction
With the rapid advancement of the powerful learning abilities of Deep Neural Networks (DNNs) [1, 2], the trustworthiness of DNNs has attracted considerable attention in recent years [3]. The training and test samples of DNNs are viewed as samples from some In-Distribution (InD) $\mathbb{P}_{\rm in}$, while the samples from other data sets are regarded as coming from a different distribution $\mathbb{P}_{\rm out}$ and are named as out-distribution data. DNNs trained on InD data output unreliable results on OoD data. Detecting whether a new sample is from $\mathbb{P}_{\rm in}$ or $\mathbb{P}_{\rm out}$ has thus been a valuable research topic of trustworthy deep learning, known as OoD detection [4].
@@
Existing OoD detection methods exploit DNNs from different aspects, e.g., logits [5], gradients [6] and features [7, 8], to unveil the disparities between InD and OoD data. In this work, we address the OoD detection from the perspective of utilizing the feature spaces learned by the backbone of DNNs. To be specific, given a DNN $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, $f$ takes $\boldsymbol{x}\in\mathbb{R}^{d}$ as inputs and learns penultimate layer features $\boldsymbol{z}\in\mathbb{R}^{m}$ before the last linear layer. Guan et al. [8] investigated the Principal Component Analysis (PCA)-based reconstruction error in the $\boldsymbol{z}$-space as the OoD detection score. That is, PCA is executed on the penultimate features $\boldsymbol{z}$ of InD training samples and learns a linear subspace spanned by the principal components from $\boldsymbol{z}$. Then, the reconstruction error measures the Euclidean distance between the original feature $\boldsymbol{z}$ and its reconstructed counterpart, which is computed by projecting $\boldsymbol{z}$ to the linear subspace and projecting it back. For good OoD detection performances, one expects that the InD features $\boldsymbol{z}$ are compactly allocated along the linear principal components with high variances for capturing most of the informative patterns of InD, which cannot be attained with the OoD features.
(a) Feature visualization of $\boldsymbol{z}$ and PCA reconstruction errors.
(b) Feature visualization of $\Phi(\boldsymbol{z})$ and KPCA reconstruction errors.
@@
Figure 1: The t-SNE [9] visualization on the primal features $\boldsymbol{z}$ (left) and the mapped features $\Phi(\boldsymbol{z})$ (right). Our KPCA detector alleviates the linear inseparability between InD and OoD data in the primal $\boldsymbol{z}$-space via an explicit feature mapping $\Phi$, and thus substantially improves the OoD detection performance, illustrated by the much more distinguishable reconstruction errors.
However, it has been observed in [8] that reconstruction errors alone cannot distinctively differentiate OoD data from InD data, leading to poor detection performance of PCA in the $\boldsymbol{z}$-space. The reasons behind lie in the intrinsic requisite for such PCA-based reconstruction errors to work: InD and OoD features should be linearly separable. Nevertheless, our observation in Fig.1(a) reveals linearly inseparable patterns among InD and OoD features in $\boldsymbol{z}$-space, which make the principal subspace extracted by PCA fail to contain informative components that could help separate OoD data from InD data.
To address this issue, we introduce non-linearity via Kernel PCA (KPCA) in $\boldsymbol{z}$-space for effective OoD detection. KPCA has long been a powerful technique in learning the non-linear patterns of data [10]. By deploying KPCA, a non-linear feature mapping $\Phi$ is imposed on the $\boldsymbol{z}$-space in this setup, so that the linear inseparability can be alleviated in the mapped $\Phi(\boldsymbol{z})$-space. KPCA is generally conducted through a kernel function $k$ induced by the feature mapping, i.e., $k(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=\langle\Phi(\boldsymbol{z}_{1}),\Phi(\boldsymbol{z}_{2})\rangle$, skipping to explicitly calculate the feature mapping $\Phi$. Regarding the OoD detection task, the key to KPCA depends on finding an appropriate $k(\cdot,\cdot)$ or $\Phi$ that introduces non-linearity to $\boldsymbol{z}$-space for promoting better separability between InD and OoD data in non-linear subspace.
In this work, we propose to leverage KPCA for OoD detection in a way of exploring explicit non-linear feature mappings $\Phi$ on penultimate features $\boldsymbol{z}$ from DNNs. To better understand the non-linear patterns in InD and OoD features, we take a kernel perspective on an existing OoD detector [7] which searches $k$-th Nearest Neighbors (KNN) on the $\ell_{2}$ normalized features $\boldsymbol{z}$. We underscore that the crucial $\ell_{2}$ normalization on $\boldsymbol{z}$ in KNN exactly suggests a feature mapping w.r.t. a cosine kernel. Motivated by this observation, we further devise two effective feature mappings $\Phi$ w.r.t. two kernels $k$ that well characterize the non-linearity in InD and OoD features. As a result, the mapped $\Phi(\boldsymbol{z})$-space enables PCA to extract an informative subspace composed of principal components that benefit the separation between InD and OoD features. As shown in Fig.1, the proposed feature mapping $\Phi$ significantly alleviates the linear inseparability of InD and OoD features in $\boldsymbol{z}$-space, leading to the substantially improved distinguishable KPCA reconstruction errors.
Our KPCA via explicit feature mappings differs from classic KPCA where kernel functions are adopted. In this work, we supplement our method with its implementation via kernel functions, and illustrate the advantageous effectiveness and efficiency of explicit feature mappings in OoD detection. Besides, extensive empirical results on multiple OoD data sets and network structures demonstrate compelling advantages of our method: (i) the efficient $\mathcal{O}(1)$ time complexity in inference over the $\mathcal{O}(N_{\rm tr})$ of KNN and KPCA using kernel functions, as the latter two methods both require that a new sample has to iterate all the $N_{\rm tr}$ training samples for nearest neighbor searching or to compute the kernel function values, respectively; (ii) the effectiveness in detecting OoD samples with new state-of-the-art (SOTA) performance over PCA reconstruction errors [8], which are still calculated in $\boldsymbol{z}$-space without taking the non-linearity in InD and OoD features into account. The contributions of this work are summarized as follows:
  • We introduce KPCA with explicit non-linear feature mappings on penultimate features from DNNs for effective OoD detection.
@@
  • Two feature mappings w.r.t. two non-linear kernels are devised to capture the non-linearity in InD and OoD, and lead to distinguishable KPCA reconstruction errors for OoD detection with complexity $\mathcal{O}(1)$ in inference.
@@
  • Analytical and experimental comparisons indicate the superiority of our KPCA detector with SOTA detection performance and remarkably reduced time complexity.
@@
In the remainder of this work, related works are outlined in Sec.2. Section 3 delves into details of our KPCA method. Experiments and analysis are presented in Sec.4 and Sec.5. Conclusions are presented in Sec.6.
@@ 2.1 Out-of-distribution Detection
Given a new sample $\boldsymbol{\hat{x}}$, the scoring function $S(\cdot)$ outputs a score $S(\boldsymbol{\hat{x}})$. If $S(\boldsymbol{\hat{x}})$ is greater than a threshold $s$, then the decision function $D(\cdot)$ would view $\boldsymbol{\hat{x}}$ as an in-distribution sample, and vice versa. The key to successfully detecting OoD samples is a well-designed scoring function. Existing OoD detectors adopt different outputs of DNNs to design justified scores to measure the disparity between InD and OoD data. A brief outline on representative OoD detectors of distinct categories is summarized below.
Logits-based detectors exploit the abnormal responses reflected in the predictive logits or probabilities from DNNs to detect out-distribution data. Typical methods adopt either the maximum probability [11, 12] or the energy function on logits [5] as the detection score.
Gradients-based methods utilize differences on gradients w.r.t. InD and OoD data for OoD detection. Gradient norms [6] or low-dimensional representations [13] are studied to devise the detection score.
Features-based detectors try to capture the feature information causing over-confidence of OoD predictions in different ways. Feature clipping [14, 15], feature distances [16, 7], feature norms [17], rank-1 features [18], feature subspace [8], etc., have been explored and shown powerful detection abilities.
Aside from methods above, other existing OoD detectors cover the training regularization [19], the ensemble technique [20] and theoretical understandings [21, 22]. Refer to the Appendix for more detailed descriptions on the compared methods in experiments [23, 24, 25, 26].
@@ 2.2 Random Fourier Features
A concise description is firstly given on the Random Fourier Features (RFFs) [27], which will be adopted in our method later. RFFs are proposed to alleviate the heavy computation cost in large-scale kernel machines. In kernel methods, an $N\times N$ kernel matrix w.r.t. $N$ samples requires $\mathcal{O}(N^{2})$ kernel manipulations, $\mathcal{O}(N^{2})$ space complexity and $\mathcal{O}(N^{3})$ time complexity to calculate the inverse of the kernel matrix. The memory and time expenses of kernel methods are even exacerbated especially for a large data size $N$. Therefore, RFFs are introduced by building an explicit feature mapping to directly approximate the kernel function for efficient kernel machines on large-scale data.
The theory foundation of RFFs is based on the Bochner’s theorem [28]: A continuous and shift-invariant kernel $k(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=k(\boldsymbol{z}_{1}-\boldsymbol{z}_{2})$ on $\mathbb{R}^{m}$ is positive definite if and only if $k(\boldsymbol{\delta})$ is the Fourier transform of a non-negative measure. 27 then derived an explicit feature mapping $\phi_{\rm RFF}$:
$\phi_{\rm RFF}(\boldsymbol{z})$ $\triangleq\sqrt{\frac{2}{M}}\left[\phi_{1}(\boldsymbol{z}),\cdots,\phi_{M}(\boldsymbol{z})\right],$ (2)
$\phi_{i}(\boldsymbol{z})$ $=\cos\ (\boldsymbol{z}^{\top}\boldsymbol{\omega}_{i}+u_{i}),i=1,\cdots,M,$
where $\boldsymbol{\omega}_{1},\cdots,\boldsymbol{\omega}_{M}\in\mathbb{R}^{m}$ are i.i.d. sampled from the Fourier transform of $k(\boldsymbol{\delta})$, and $u_{1},\cdots,u_{M}\in\mathbb{R}$ are i.i.d. sampled from the uniform distribution $\mathcal{U}(0,2\pi)$. For example, the Fourier transform of a Gaussian kernel function $k_{\rm gau}=e^{-\gamma\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\|^{2}_{2}}$ is $\mathcal{N}(0,\sqrt{2\gamma}\boldsymbol{I}_{m})$. Such a feature mapping $\phi_{\rm RFF}$ satisfies $k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})\approx\phi_{\rm RFF}(\boldsymbol{z}_{1})^{\top}\phi_{\rm RFF}(\boldsymbol{z}_{2})$ and is known as the random Fourier features. Refer to [27] for a detailed convergence analysis. RFFs have been widely utilized in kernel learning [29], optimization [30], etc.

### v3

@@ Abstract
Out-of-Distribution (OoD) detection is vital for the reliability of Deep Neural Networks (DNNs). Existing works have shown the insufficiency of Principal Component Analysis (PCA) straightforwardly applied on the features of DNNs in detecting OoD data from In-Distribution (InD) data. The failure of PCA suggests that the network features residing in OoD and InD are not well separated by simply proceeding in a linear subspace, which instead can be resolved through proper non-linear mappings. In this work, we leverage the framework of Kernel PCA (KPCA) for OoD detection, and seek suitable non-linear kernels that advocate the separability between InD and OoD data in the subspace spanned by the principal components. Besides, explicit feature mappings induced from the devoted task-specific kernels are adopted so that the KPCA reconstruction error for new test samples can be efficiently obtained with large-scale data. Extensive theoretical and empirical results on multiple OoD data sets and network structures verify the superiority of our KPCA detector in efficiency and efficacy with state-of-the-art detection performance.
@@ 1 Introduction
With the rapid advancement of the powerful learning abilities of Deep Neural Networks (DNNs) [1; 2], the trustworthiness of DNNs in security-sensitive scenarios has attracted considerable attention in recent years [3]. Generally, samples from the training and test sets of DNNs are viewed as data from some In Distribution (InD) $\mathbb{P}_{\rm in}$, while samples from other data sets are regarded as coming from a different distribution $\mathbb{P}_{\rm out}$, i.e., out-of-distribution (OoD) data. In the practical deployment, DNNs trained on InD data would inevitably encounter OoD data and thus yield unreliable results with potential risks. Therefore, detecting whether a new sample is from $\mathbb{P}_{\rm in}$ or $\mathbb{P}_{\rm out}$ has been a valuable research topic of trustworthy deep learning, namely OoD detection [4].
@@
Existing OoD detection methods exploit different outputs from DNNs to unveil the hidden disparities between InD and OoD data, e.g., logits [5], gradients [6] and features [7; 8]. In this work, we address OoD detection from a perspective of utilizing the feature spaces learned by the backbone of DNNs. To be specific, given a DNN $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, $f$ takes $\boldsymbol{x}\in\mathbb{R}^{d}$ as inputs and learns penultimate layer features $\boldsymbol{z}\in\mathbb{R}^{m}$ before the last linear layer. Principal Component Analysis (PCA) is investigated in [8] to calculate the reconstruction error in the $\boldsymbol{z}$-space as the OoD detection score. That is, PCA is executed on the penultimate features of InD training samples and learns a linear subspace spanned by the principal components in the $\boldsymbol{z}$-space. Then, given features $\boldsymbol{\hat{z}}$ of an unknown sample $\boldsymbol{\hat{x}}$, one can obtain the reconstructed counterpart of $\boldsymbol{\hat{z}}$ by projecting $\boldsymbol{\hat{z}}$ to the linear subspace and re-projecting it back. The reconstruction error is computed as the Euclidean distance between $\boldsymbol{\hat{z}}$ and the reconstructed counterpart. In [9], similar ideas are adopted with energy-based models using an auto-encoder structure, where the neural networks are trained from scratch and the reconstruction is conducted in the decoder. For good OoD detection performance, it is expected that InD features are compactly allocated along the linear principal components with high variances for capturing most of the informative patterns of InD data, leading to small reconstruction errors, while OoD features are not supposed to be well matched with the learned subspace, causing large reconstruction errors.
(a) T-SNE of $\boldsymbol{z}$ and PCA reconstruction errors.
(b) T-SNE of $\Phi(\boldsymbol{z})$ and KPCA reconstruction errors.
@@
Figure 1: The t-SNE [10] visualization on the original features $\boldsymbol{z}$ (left) and the mapped features $\Phi(\boldsymbol{z})$ (right). Our KPCA detector alleviates the linearly inseparability between InD and OoD features in the original $\boldsymbol{z}$-space via an explicit feature mapping $\Phi$, and thus substantially improves the OoD detection performance, illustrated by the much more distinguishable reconstruction errors.
However, it has been empirically observed in [8] that such PCA reconstruction errors alone cannot distinctively differentiate OoD data from InD data, leading to poor detection performance of PCA in the $\boldsymbol{z}$-space. Nevertheless, [8] did not take further explorations on the reasons behind, and instead proposed a practical fusion trick to boost PCA by multiplying with other existing powerful detection scores. Therefore, in this work, more in-depth analyses are undertaken to improve the limitations of PCA for OoD detection with insightful understandings on the distribution of InD and OoD features. It is widely acknowledged that PCA falls inferior in dealing with those linearly-inseparable data, which inspires us to explore the non-linearity existing in the $\boldsymbol{z}$-space of InD and OoD features under the help of the celebrated Kernel PCA (KPCA).
KPCA has long been a powerful technique in learning the non-linear patterns of data [11; 12]. By deploying KPCA, a non-linear feature mapping $\Phi$ is imposed on the $\boldsymbol{z}$-space in this setup, so that the linear inseparability can be alleviated in the mapped $\Phi(\boldsymbol{z})$-space. KPCA is generally conducted through a kernel function $k$ induced by the feature mapping, i.e., $k(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=\langle\Phi(\boldsymbol{z}_{1}),\Phi(\boldsymbol{z}_{2})\rangle$, in order to avoid explicit calculations in the mapped $\Phi(\boldsymbol{z})$-space. In most cases, researchers have no prior on the unknown non-linear data distribution, e.g., the non-linearity in the $\boldsymbol{z}$-space of InD and OoD features. Therefore, finding an appropriate $k$ or $\Phi$ that well adapts the data always remains a non-trivial issue for KPCA. For example, the mostly common Gaussian kernel is shown to be unable to separate the InD and OoD features, and leads to terrible OoD detection performance, see details in Appendix C. In addition, KPCA also faces the challenge of calculating and storing the $N_{\rm tr}\times N_{\rm tr}$ kernel matrix on millions of training samples with a very large size $N_{\rm tr}$, which significantly hinders its application in practical tasks with a huge amount of data.
The proposed KPCA detection method well addresses the aforementioned issues of KPCA for OoD detection. On the one hand, to better understand the non-linear patterns in InD and OoD features, we take a kernel perspective on an existing OoD detector [7], which searches $k$-th Nearest Neighbors (KNN) on the $\ell_{2}$ normalized features $\boldsymbol{z}$. By decoupling and analyzing key components of the KNN method, we acquire two effective kernels, a cosine kernel and a cosine-Gaussian kernel, for our KPCA detector to promote the linear separability between InD and OoD features in the subspace of the principal components, leading to substantially improved distinguishable KPCA reconstruction errors, as shown in Figure 1. On the other hand, for a computationally-friendly implementation, two explicit feature mappings $\Phi$ induced from the cosine and cosine-Gaussian kernels are executed on the original features $\boldsymbol{z}$, followed by PCA in the mapped $\Phi(\boldsymbol{z})$-space to obtain the reconstruction errors without calculations on the kernel matrix. Specifically, the celebrated Random Fourier Features (RFFs) [13] are introduced to approximate the Gaussian kernel, allowing an $\mathcal{O}(M)$ computation complexity in inference, which significantly outperforms the $\mathcal{O}(N_{\rm tr})$ computation complexity of the KNN [7] and the kernel-matrix-based KPCA ($M$ is the number of RFFs and $M\ll N_{\rm tr}$).
Extensive experiments verify the effectiveness of the devoted two kernels, which we hope could bring inspirations for the research community in exploring the non-linearity in InD and OoD data from a kernel perspective. For example, the two kernels can even serve as a beneficial prior on advocating learning more and stronger kernels for OoD detection. In addition, we supplement our method with its implementation via the kernel matrix, and illustrate the advantageous effectiveness and efficiency of explicit feature mappings in Section 6. The contributions of this work are summarized below:
  • To the best of our knowledge, this is the first work that explores suitable kernels to seize the non-linearity in InD and OoD features in the post-hoc stage on well-trained DNNs.
@@
  • Two task-specific kernels are carefully devised for OoD detection. Particularly, two explicit feature mappings induced from the kernels are adopted for the KPCA detector, and lead to separable KPCA reconstruction errors with significantly-reduced complexity in inference.
@@
  • Theoretical and experimental comparisons indicate the effects of kernels in our KPCA detector with SOTA detection performance and remarkably reduced time complexity.
@@
In the remainder of this work, related works and research backgrounds are outlined in Section 2 and Section 3, respectively. Section 4 delves into details of the proposed KPCA detector. Comparison experiments with prevailing OoD detection methods and KPCA via kernel functions are presented in Section 5 and Section 6, respectively. Conclusions and limitations are drawn in Section 7.
@@ 2 Related work
The scoring function $S(\cdot)$ assigns a score $S(\boldsymbol{\hat{x}})$ for a new sample $\boldsymbol{\hat{x}}$. If $S(\boldsymbol{\hat{x}})$ is greater than a threshold $s$, the decision function $D(\cdot)$ would view $\boldsymbol{\hat{x}}$ as an InD sample, and vice versa. The key to effectually detecting OoD samples is a well-designed scoring function. Existing OoD detectors adopt different outputs from DNNs and design justified scores to measure the disparity between InD and OoD data.
Logits-based detectors exploit the abnormal responses reflected in the predictive logits or probabilities from DNNs to detect OoD data. Typical methods adopt either the maximum logits [14] or probability [15; 16] or the energy function on logits [5] as the detection score.
Gradients-based methods investigate differences on gradients w.r.t InD and OoD data for OoD detection. For example, gradient norms [6] or low-dimensional representations [17] are studied to devise the detection score.
Features-based detectors try to capture the feature information causing over-confidence of OoD predictions in different ways. Feature clipping [18; 19; 20; 21; 22], feature distances [23; 7; 24], feature norms [25], rank-1 features [26], feature subspace [8], etc., have been explored with excellent performance.
Aside from methods above, other existing OoD detectors cover the training regularization [27; 28], the ensemble technique [29] and theoretical understandings [30; 31]. Refer to Appendix A for more detailed descriptions on the compared methods in experiments [32; 33; 34; 35].
3 Background
3.1 PCA for OoD detection
$\boldsymbol{\Sigma}=\sum_{i=1}^{N_{\rm tr}}\left(\boldsymbol{z}_{i}-\boldsymbol{\mu}\right)\left(\boldsymbol{z}_{i}-\boldsymbol{\mu}\right)^{\top},\quad\boldsymbol{\mu}=\frac{1}{N_{\rm tr}}\sum_{i=1}^{N_{\rm tr}}\boldsymbol{z}_{i}.$ (2)
Through the eigendecomposition $\boldsymbol{\Sigma}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}$, the dimensionality reduction matrix $\boldsymbol{U}_{q}\in\mathbb{R}^{m\times q}$ is obtained by taking the first $q$ columns of the eigenvector matrix $\boldsymbol{U}$ w.r.t the top-$q$ largest eigenvalues.
In inference, given a new sample $\boldsymbol{\hat{x}}\in\mathbb{R}^{d}$ and its feature $\boldsymbol{\hat{z}}\in\mathbb{R}^{m}$ from the DNN $f$, the reconstruction error is computed as:
By projecting centralized $(\boldsymbol{\hat{z}}-\boldsymbol{\mu})$ to the $\boldsymbol{U}_{q}$-subspace and re-projecting back, we can obtain the reconstructed features $\boldsymbol{U}_{q}\boldsymbol{U}_{q}^{\top}(\boldsymbol{\hat{z}}-\boldsymbol{\mu})$ and the reconstruction error $e(\boldsymbol{\hat{x}})$, which then can be set as the OoD detection score: $S(\boldsymbol{\hat{x}})=-e(\boldsymbol{\hat{x}})$. An ideal case is that $\boldsymbol{U}_{q}$ contains informative principal components of InD data and causes projections of OoD data far away from that of InD data, leading to separable reconstruction errors between OoD and InD data.
@@ 3.2 Random Fourier features
A concise description is firstly given on the Random Fourier Features (RFFs) [13], which will be adopted in our method later. RFFs are proposed to approximate the kernel function so as to alleviate the heavy computation cost in large-scale kernel machines. In kernel methods, an $N\times N$ kernel matrix w.r.t $N$ samples requires $\mathcal{O}(N^{2})$ kernel manipulations, $\mathcal{O}(N^{2})$ space complexity and $\mathcal{O}(N^{3})$ time complexity to calculate the inverse of the kernel matrix, which leads to overwhelmed computation costs for a large data size $N$. Therefore, RFFs are introduced by building an explicit feature mapping to directly approximate the kernel function for efficient kernel machines on large-scale data.
RFFs are built on the Bochner’s theorem [36]: A continuous and shift-invariant kernel $k(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=k(\boldsymbol{z}_{1}-\boldsymbol{z}_{2})$ on $\mathbb{R}^{m}$ is positive definite if and only if $k(\cdot)$ is the Fourier transform of a non-negative measure. An explicit feature mapping $\phi_{\rm RFF}$ induced from the kernel $k$ is derived in [13]:
$\phi_{\rm RFF}(\boldsymbol{z})\triangleq\sqrt{\frac{2}{M}}\left[\phi_{1}(\boldsymbol{z}),\cdots,\phi_{M}(\boldsymbol{z})\right],\quad\phi_{i}(\boldsymbol{z})=\cos\ (\boldsymbol{z}^{\top}\boldsymbol{\omega}_{i}+u_{i}),i=1,\cdots,M,$ (4)
where $\boldsymbol{\omega}_{1},\cdots,\boldsymbol{\omega}_{M}\in\mathbb{R}^{m}$ are i.i.d. sampled from the Fourier transform of $k(\cdot)$, and $u_{1},\cdots,u_{M}\in\mathbb{R}$ are i.i.d. sampled from a uniform distribution $\mathcal{U}(0,2\pi)$. For example, the sampling distribution for $\boldsymbol{\omega}_{i}$ of a Gaussian kernel function $k_{\rm gau}=e^{-\gamma\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\|^{2}_{2}}$ is $\boldsymbol{\omega}\sim\mathcal{N}(0,\sqrt{2\gamma}\boldsymbol{I}_{m})$. Such a feature mapping $\phi_{\rm RFF}$ satisfies $k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})\approx\phi_{\rm RFF}(\boldsymbol{z}_{1})^{\top}\phi_{\rm RFF}(\boldsymbol{z}_{2})$ and is known as the random Fourier features (RFFs) mapping. Refer to [13] for a detailed convergence analysis. RFFs have been widely utilized in kernel learning [37], optimization [38], etc.